OpenStudy (anonymous):

a side of a square is 16 inches. The midpoints of its sides are joined to form an inscribed square. Another square is drawn in such a way that its vertices would lie also at the midpoints of the sides of the second square. This process is continued infinitely. Find the sum of the areas of these infinite squares.

OpenStudy (anonymous):

help please

OpenStudy (akashdeepdeb):

Do you know what geometric progressions are?

OpenStudy (anonymous):

no sir

OpenStudy (anonymous):

this uses geometric progression

OpenStudy (anonymous):

what is geometric progression?

OpenStudy (anonymous):

Geometric progression is a vast topic but u can check it on http://en.wikipedia.org/wiki/Geometric_progression and see how much u have and i will type the answer to your question.

OpenStudy (anonymous):

It appears that this forms a geometric series with a ratio of .5 The formula for an infinite geometric series is: S[infinity] = a[1]/(1-r) where: a[1] = the first area = 16*16 = 256 r = the ratio = .5 This formula becomes: S[infinity] = 256/(1-.5) = 256/.5 = 512 Sum should equal 512. Here's how it was derived. The first square is 16 * 16 = 256 square inches. The first inscribed square is 11.3137085 * 11.3137085 = 128 square inches. The second inscribed square is 8 * 8 = 64 square inches. The third inscribed square is 5.656854249 * 5.656854249 = 32 square inches. The fourth inscribed square is 4 * 4 = 16 square inches. Each succeeding inscribed square is 1/2 * the area of the one above it. Adding them up, you get: 256 + 128 + 64 + 32 + 16 + 8 + 4 + 2 + 1 = 511 + ever decreasing fractions which will converge on 512 as the number of iterations increases towards infinity. The formula for finding the side of the inscribed square is the side of the original square times cos(45). original is 16 16 * cos(45) = 11.3137085 11.3137085 * cos(45) = 8 8 * cos(45) = ......... Check this Picture of Inscribed Square to see how the length of each side of the inscribed square was derived. Each side of the inscribed square formed a triangle with the corner of the original square. The side of the inscribed square was the hypotenuse of the triangle. The length of the hypotenuse of each triangle was equal to half the length of the original side * cos(45). This became .707106781 * the length of the original side. Original Side was 16 Inscribed square side became 11.3137.... Next Inscribed square side became .707 * 11.313 = 8 etc.

OpenStudy (anonymous):

can i see the figure of the problem?

OpenStudy (akashdeepdeb):